A) What is the magnitude of the electric force between charges of 0.29 C and 0.15...

A)

What is the magnitude of the electric force between charges of 0.29 C and 0.15 C at a separation of 0.85 m ? k=8.99×109N⋅m2/C2.

Express your answer to two significant figures and include the appropriate units.

B)

If the separation between the charges is increased, does the magnitude of the force increase, decrease, or stay the same?

If the separation between the charges is increased, does the magnitude of the force increase, decrease, or stay the same?

stay the same
decrease
increase
Answer #1

A) Magnitude of electric force is given by, F = kq1q2/r2

k = 1/4o = 8.99*109 N-m2/C2

q1 = 0.29 C; q2= 0.15 C; r = 0.85 m

F = 8.99*109*0.29*0.15/(0.852)

F = 5.412*108 N

B) From the formula given above. we can see that as r increases F decreases